why magnetic field due to infinitely  long wire and finite long wire comes out to be same using amperes law?

Grade:12

1 Answers

Eshan
5 years ago
Dear student,

If this is so, you are making a mistake while applying the Ampere’s law. Ampere’s law only states that line integral of magnetic field is equal to\mu_0 I, which is always true.

Hence\int\vec{B}.d\vec{l}=\mu_0Iis always true.

However if you expanded it to\int\vec{B}.d\vec{l}=B(2\pi r), you have assumed yourself that magnitude of magnetic field is constant as well as ‘along’ the circular path, which is a wrong assumption. Hence by using you cannot proceed further from\int\vec{B}.d\vec{l}=\mu_0 I, and you would have to use biot-savart’s law.